OpenStudy (anonymous):

f(x)=sq rt (4-7x) can some one please explain how to find the range and domian of this equation

OpenStudy (jdoe0001):

domain will be the values "x" can take range will be the values "y" can take, keep in mind that "y" values depend on whatever value "x" turns into

OpenStudy (anonymous):

so the domian wi be (-infiny,+infinty) since its doesnt equal zero?

OpenStudy (jdoe0001):

so as far as roots with even radical, that is square root in this case restrictions inside radical pertain to the issue that the radical amount inside CANNOT BE NEGATIVE so any values "x" will consider in taking, will have to take that into consideration so, what values can "x" take without making the radical negative? that is, what "x" can be without making 4-7x a negative figure?

OpenStudy (jdoe0001):

well, let's see on that, let's use the value x = 1 4-7x => 4-7(1) => -3 well, that's negative and \(\bf \sqrt{-3} \implies i\sqrt{3}\) which is an imaginary number so 1 for "x" is not good, since it makes the radical value negative

OpenStudy (anonymous):

oh okay give me a second to figure it out and ask you if its right

OpenStudy (jdoe0001):

notice that \(\bf \sqrt{0} = 0\) so that's ok, it's just when you get into the negative when the issue turns to imaginary

OpenStudy (anonymous):

Okay so since you start of at zero and can positive its (-infiy,0]

OpenStudy (anonymous):

can't*

OpenStudy (jdoe0001):

hmm, lemme give you a few values there \(\bf f(x)=\sqrt{4-7x}\\ x= 1, \qquad \sqrt{4-7(1)} \implies \sqrt{-3}\qquad \textit{no good}\\ x= 0, \qquad \sqrt{4-7(0)} \implies \sqrt{4}\qquad \textit{good}\\ x = \cfrac{4}{7} \qquad \sqrt{4-7\left(\frac{4}{7} \right)} \implies \sqrt{0}\qquad \textit{good}\\ \textit{if "x" goes above }\cfrac{4}{7} \textit{then you get negative values for the radical}\)

OpenStudy (jdoe0001):

so the values "x" can safely take without making the radical negative are, any negative value up to \(\bf \cfrac{4}{7} \implies (-\infty, \cfrac{4}{7}]\)

OpenStudy (anonymous):

So this would be the range correct? and this only accounts for problems that have square roots.

OpenStudy (jdoe0001):

the range consequently due to that restriction to "x", will be that yes

OpenStudy (jdoe0001):

as far as "x", x can take anything

OpenStudy (jdoe0001):

this restriction accounts to roots that have an EVEN RADICAL that is \(\bf \Large \sqrt[4]{x}\qquad \sqrt[6]{x}\qquad \sqrt[20]{x} ...\)

OpenStudy (jdoe0001):

on an ODD RADICAL that doesn't happen, an example of that will be \(\bf \sqrt[3]{-27} \implies \sqrt[3]{-3\times -3 \times -3} \implies \sqrt[3]{(-3)^3}\implies -3\)

OpenStudy (jdoe0001):

a negative number can have an ODD root, thus it doesn't happen with ODD RADICALS, just with EVEN ones

OpenStudy (anonymous):

Okay so the range is (-infy,4/7] and the domain is [0,infy)

OpenStudy (jdoe0001):

the domain will be \(\bf (-\infty, \quad +\infty)\)

OpenStudy (jdoe0001):

so, "x" can take anything, but "y" will be restricted to only the positive values in the radical

OpenStudy (jdoe0001):

hmmm, not sure if I'm confusing... maybe I'm ... heheh, the domain and range,

OpenStudy (anonymous):

Im a bit confused im trying to learn this out of pauls note introduction ex 4 letter b and it explains it differently

OpenStudy (jdoe0001):

I think I got them backwards, the restrictions to "x" are to the domain

OpenStudy (jdoe0001):

as far as the range, well, oddly enough because an even root can spit out negative and positive roots, will be \(\bf (-\infty, \quad +\infty)\)

OpenStudy (jdoe0001):

why, well, because \(\Large y = \pm \sqrt{4-7x}\) so any value the radical becomes, the root itself will give a positive and negative account of it

OpenStudy (jdoe0001):

am I making any sense? hehe

OpenStudy (anonymous):

Give me a sec im read everything that you explained and he has and let you know if i get it. lol

OpenStudy (jdoe0001):

\(\bf y = \pm \sqrt{4-7x}\\ x = -100 \qquad y = \pm \sqrt{74} \implies \pm 8.6\\ x = -1,000,000 \qquad y = \pm \sqrt{1,000,004} \implies \pm 1000.002\)

OpenStudy (anonymous):

do you do tuyoring by any chance?

OpenStudy (jdoe0001):

so the \(\bf \Large \pm \) part of the root, will give all real numbers

OpenStudy (jdoe0001):

could, but no

OpenStudy (anonymous):

I under stand now. Thanks a bunch. Have a great day